OpenStudy (anonymous):

Find fog if f(x)= 2x-1 and g(x)= x+2

OpenStudy (anonymous):

f o g = f of g any ideas how to start?

OpenStudy (anonymous):

hai sugga daddi

OpenStudy (anonymous):

\[f o g = f(g(x))\]

OpenStudy (anonymous):

where's my kisses bad

OpenStudy (anonymous):

Find fog if f(x)= 2x-1 and g(x)= x+2 A) 2x-3 B) 2x+1 C) 2x-1 D) 2x+3

OpenStudy (anonymous):

bae

OpenStudy (natsent06):

@iambatman is it 16%

OpenStudy (anonymous):

Yes, lissa we don't need the options, but can you tell us what to do next with the information I gave you?